The area of no return at one intersection was calculated to be 59.0 m. What is the minimum length of time the traffic light should be yellow so a vehicle just entering the area of no return can clear the area while maintaining a speed of 70.0 km/h?


11.9 s


9.10 s


6.07 s


3.03 s


1
Expert's answer
2021-11-11T12:04:19-0500

"v = 70\frac{km}{h}"

"s = 59.0m"

"v =70* \frac{1000}{3600}\approx 19.44\frac{m}{s}"

"s =v*t"

"t = \frac{s}{v}= \frac{59}{19.44}\approx 3.03s"


"\text{Answer: }3.03\ s"